# Spurious Privacy Leakage in Neural Networks

### v1

@@ Appendix B Spurious privacy leakage
We report additional technical details related to Section 3 and include additional results: formalizing the feature similarity and complexity definitions, comparing different membership inference attacks on spurious data, demonstrating how memorization of spurious data causes higher privacy leakage.
Hyperparameters. For Section 3, we apply grid search to find the best hyperparameters for each dataset. For Waterbirds and CelebA we search the learning rate between [1e-3, 1e-4] and weight decay [1e-1, 1e-2, 1e-3]. For FMoW the learning rate [1e-3, 3e-3, 1e-4, 3e-4], weight decay [1e-1, 1e-2, 1e-3], and epochs [20, 30, 40]. For MultiNLI the learning rate [1e-5, 3e-5], weight decay [1e-5, 1e-4]. For CivilComments the learning rate [1e-5, 1e-6], weight decay [1e-3, 1e-4]. The best hyperparameters are reported at Table 3.
@@ B.1 Measuring task complexity
We formalize the measures used in Section 3.2 to quantity the feature similarity and feature complexity between different groups of data.
@@ B.2 Membership inference attacks comparison
Most of the previous MIAs are limited by the assumption that all the samples have the same level of importance (or hardness) (Yeom et al., 2018; Shokri et al., 2017), which is incorrect since natural data follow a long-tail distribution (Feldman, 2020). We compare three different state-of-the-art MIAs and show that the phenomenon of spurious privacy leakage exists regardless of the attack used. We use two different versions of LiRA (Carlini et al., 2022), online and offline, and TrajMIA (Liu et al., 2022a). The results in Table 4 show that all the methods successfully reveal the disparity on Waterbirds, and LiRA online is the strongest attack on vulnerable groups.
@@ B.3 Memorization score of spurious groups
Feldman (2020) introduced the notion of label memorization (Definition B.3) as the difference in the label of a model trained with or without $\bm{x}$. We use the models from the LiRA algorithm from Section 3.1 to approximate the memorization score. Carlini et al. (2022) proposed the privacy score $d=|\mu_{\text{in}}-\mu_{\text{out}}|/(\sigma_{\text{in}}+\sigma_{\text{out}})$ to measure the difference between the loss distributions coming from IN and OUT shadow models of LiRA. Note that both mem(.) and $d$ measure the difference between two probability distributions conditioned on $D$ and $D\setminus\{i\}$ but with a different level of granularity; label memorization is coarser than $d$ and collapses the whole distributions to a single scalar, the probability of outputting the correct label.
@@
We compute $d$ for each data point and use a Gaussian kernel density estimator to fit each group. The results in Figure 7 show the estimated frequency of the memorization score for the whole dataset divided per group. We observe that the spurious groups have, on average, higher memorization scores compared to non-spurious groups (except for FMoW as in Figure 1). The increase can be attributed to the presence of spurious features, which turn typical examples into atypical ones that the model has to memorize. A higher memorization score is known to be linked to a higher vulnerability under privacy attacks (Feldman, 2020), which matches what we observed previously.
@@ Appendix C Robust training
We report additional technical details related to Section 4.1 and include an additional result analyzing the privacy side effect of choosing L2 vs L1 regularization in DFR.
Hyperparameters. We use the same hyperparameters as in Table 3. Robust training DRO requires an extra hyperparameter C. For Waterbirds and CelebA we tune C within [0, 1, 2, 3, 4], for FMoW [0, 1, 2, 4, 8, 16], and for MultiNLI and CivilComments [0, 1, 2, 4, 8, 16]. For DFR, we do not use the validation set for retraining but use a group-balanced subset sampled from the training set. This allows a fairer comparison with other methods by not exploiting additional data, and it is also necessary for a fair privacy analysis since adding extra data invalidates the membership inference comparison.
@@
Data Model Train Acc. ($\uparrow$) Test Acc. ($\uparrow$) $\Delta$Acc. ($\downarrow$) Train WGA ($\uparrow$) Test WGA ($\uparrow$) $\Delta$WGA ($\downarrow$)
@@
Experiment setup. For the LiRA attack, we train 32 ERM shadow models for Waterbirds and CelebA and 16 ERM shadow models for FMoW, MultiNLI, and CivilComments. We also train 5 DRO and DFR models for all the datasets. We use the online version of LiRA with a fixed variance for all the attacks to audit the privacy level. Table 1 reports the mean and standard error of using the ERM trained shadow models to attack 32 target models for each training type of Waterbirds and CelebA, and 5 for FMoW and MultiNLI.
We found DRO to be unstable on more complex datasets such as FMoW, where it fails to improve the validation WGA even after an extensive hyperparameter grid search. While for DFR, despite its simplicity and effectiveness compared to DRO, we find that it can slightly increase the vulnerability of spurious groups. However, by simply changing DFR’s regularization from L1 to L2 norm, we achieve an accuracy-privacy tradeoff reducing the vulnerability to the same level as ERM at the cost of a lower WGA (77.67% to 73.00%) (see Section C.1).
Table 6: Comparing the attack success rate of different training methods for spurious and non-spurious groups. This it the full version of the Table 1 in the main text. The spurious groups are highlighted. *The spurious groups 1 and 2 of Waterbirds are evaluated at 1% and 3% respectively due to the limited samples. DRO fails to improve the accuracy on FMoW after an extensive grid search, therefore we omit it.
@@ C.1 DFR with L2 regularization
DFR with the L1 regularization achieves the best performance measured with WGA. The L1 regularization encourages sparsity of the last-linear layer, concentrating most of the weights to 0. Kirichenko et al. (2022) showed that using L2 regularization leads to suboptimal results in terms of WGA performance. Additionally, we find that L2 leads to an accuracy-privacy tradeoff where it slightly increases privacy protection against MIA. We compare DFR trained with L2 and the default L1 regularization, finding that L2 regularization achieves a lower 83.65% test accuracy compared to 85.96% of L1, and also a lower WGA 73.00% compared to 77.67%. However, in Table 7, we observe that by using L2 regularization, the privacy vulnerability is reduced to a similar level of ERM.
@@ Appendix D Differential privacy
We report additional experimental details for the Section 4.2.
@@ Appendix E Architecture influences
We report additional technical details related to Section 5.
Experimet setup. For each model, we use the same computational budget by performing the grid search with the learning rate [1e-1, 1e-2, 1e-3, 1e-4] and the weight decay in [1e-1, 1e-2, 1e-3] and choose the best performing based on the validation set. For privacy analysis, we avoid overfitting by tuning the number of training epochs to stop the training before reaching 100% training accuracy (Carlini et al., 2022). The best hyperparameters are reported at Table 8.
Table 8: Final hyperparameters used for training the various architecture in Section 5.
@@ Appendix F Compute resources
All the experiments are run on our internal cluster with the GPU Tesla V100 16GB/32GB of memory. We give an estimate of the amount of compute required for each experiment. For Section 3, we trained 96 shadow models for Waterbirds and CelebA, and 48 for FMoW, MultiNLI, and CivilComments which took ~600 hours of computing. We trained 16 shadow models for FMoW4 and FMoW16 which took another ~100 hours. For Section 5, we trained in total 128 shadow models on Waterbirds and CelebA averaging around ~100 hours. Lastly for Section 4.2, we trained 32 ConvNext-t shadow models and 5 target models for about ~100 hours. The full research required additional computing for hyperparameter grid searches, in particular for differential privacy training which is known to be difficult.

### v2

@@ Appendix B Spurious privacy leakage
We report additional technical details related to Section˜3 and include additional results: formalizing the feature similarity and complexity definitions, comparing different membership inference attacks on spurious data, demonstrating how memorization of spurious data causes higher privacy leakage.
Hyperparameters. For Section˜3, we apply grid search to find the best hyperparameters for each dataset. For Waterbirds and CelebA we search the learning rate between [1e-3, 1e-4] and weight decay [1e-1, 1e-2, 1e-3]. For FMoW the learning rate [1e-3, 3e-3, 1e-4, 3e-4], weight decay [1e-1, 1e-2, 1e-3], and epochs [20, 30, 40]. For MultiNLI the learning rate [1e-5, 3e-5], weight decay [1e-5, 1e-4]. For CivilComments the learning rate [1e-5, 1e-6], weight decay [1e-3, 1e-4]. The best hyperparameters are reported at Table˜3.
Figure 7: Spurious privacy leakage is absent for the FMoW dataset. See the Section˜3.2.
@@ B.1 Measuring task complexity
We formalize the measures used in Section˜3.2 to quantity the feature similarity and feature complexity between different groups of data.
@@ B.2 Membership inference attacks comparison
Most of the previous MIAs are limited by the assumption that all the samples have the same level of importance (or hardness) (Yeom et al., 2018; Shokri et al., 2017), which is incorrect since natural data follow a long-tail distribution (Feldman, 2020). We compare three different state-of-the-art MIAs and show that the phenomenon of spurious privacy leakage exists regardless of the attack used. We use two different versions of LiRA (Carlini et al., 2022), online and offline, and TrajMIA (Liu et al., 2022a). The results in Table˜4 show that all the methods successfully reveal the disparity on Waterbirds, and LiRA online is the strongest attack on vulnerable groups.
@@ B.3 Memorization score of spurious groups
Feldman (2020) introduced the notion of label memorization (Definition˜B.3) as the difference in the label of a model trained with or without $\bm{x}$. We use the models from the LiRA algorithm from Section˜3.1 to approximate the memorization score. Carlini et al. (2022) proposed the privacy score $d=|\mu_{\text{in}}-\mu_{\text{out}}|/(\sigma_{\text{in}}+\sigma_{\text{out}})$ to measure the difference between the loss distributions coming from IN and OUT shadow models of LiRA. Note that both mem(.) and $d$ measure the difference between two probability distributions conditioned on $D$ and $D\setminus\{i\}$ but with a different level of granularity; label memorization is coarser than $d$ and collapses the whole distributions to a single scalar, the probability of outputting the correct label.
@@
We compute $d$ for each data point and use a Gaussian kernel density estimator to fit each group. The results in Figure˜8 show the estimated frequency of the memorization score for the whole dataset divided per group. We observe that the spurious groups have, on average, higher memorization scores compared to non-spurious groups (except for FMoW as in Figure˜1). The increase can be attributed to the presence of spurious features, which turn typical examples into atypical ones that the model has to memorize. A higher memorization score is known to be linked to a higher vulnerability under privacy attacks (Feldman, 2020), which matches what we observed previously.
@@ Appendix C Robust training
We report additional technical details related to Section˜4.1 and include an additional result analyzing the privacy side effect of choosing L2 vs L1 regularization in DFR.
Hyperparameters. We use the same hyperparameters as in Table˜3. Robust training DRO requires an extra hyperparameter C. For Waterbirds and CelebA we tune C within [0, 1, 2, 3, 4], for FMoW [0, 1, 2, 4, 8, 16], and for MultiNLI and CivilComments [0, 1, 2, 4, 8, 16]. For DFR, we do not use the validation set for retraining but use a group-balanced subset sampled from the training set. This allows a fairer comparison with other methods by not exploiting additional data, and it is also necessary for a fair privacy analysis since adding extra data invalidates the membership inference comparison.
@@
[…]
@@
Experiment setup. For the LiRA attack, we train 32 ERM shadow models for Waterbirds and CelebA and 16 ERM shadow models for FMoW, MultiNLI, and CivilComments. We also train 5 DRO and DFR models for all the datasets. We use the online version of LiRA with a fixed variance for all the attacks to audit the privacy level. Table˜1 reports the mean and standard error of using the ERM trained shadow models to attack 32 target models for each training type of Waterbirds and CelebA, and 5 for FMoW and MultiNLI.
We found DRO to be unstable on more complex datasets such as FMoW, where it fails to improve the validation WGA even after an extensive hyperparameter grid search. While for DFR, despite its simplicity and effectiveness compared to DRO, we find that it can slightly increase the vulnerability of spurious groups. However, by simply changing DFR’s regularization from L1 to L2 norm, we achieve an accuracy-privacy tradeoff reducing the vulnerability to the same level as ERM at the cost of a lower WGA (77.67% to 73.00%) (see Section˜C.1).
Table 6: Comparing the attack success rate of different training methods for spurious and non-spurious groups. This it the full version of the Table˜1 in the main text. The spurious groups are highlighted. *The spurious groups 1 and 2 of Waterbirds are evaluated at 1% and 3% respectively due to the limited samples. DRO fails to improve the accuracy on FMoW after an extensive grid search, therefore we omit it.
@@ C.1 DFR with L2 regularization
DFR with the L1 regularization achieves the best performance measured with WGA. The L1 regularization encourages sparsity of the last-linear layer, concentrating most of the weights to 0. Kirichenko et al. (2022) showed that using L2 regularization leads to suboptimal results in terms of WGA performance. Additionally, we find that L2 leads to an accuracy-privacy tradeoff where it slightly increases privacy protection against MIA. We compare DFR trained with L2 and the default L1 regularization, finding that L2 regularization achieves a lower 83.65% test accuracy compared to 85.96% of L1, and also a lower WGA 73.00% compared to 77.67%. However, in Table˜7, we observe that by using L2 regularization, the privacy vulnerability is reduced to a similar level of ERM.
@@ Appendix D Differential privacy
We report additional experimental details for the Section˜4.2.
@@ Appendix E Architecture influences
We report additional technical details related to Section˜5.
Experimet setup. For each model, we use the same computational budget by performing the grid search with the learning rate [1e-1, 1e-2, 1e-3, 1e-4] and the weight decay in [1e-1, 1e-2, 1e-3] and choose the best performing based on the validation set. For privacy analysis, we avoid overfitting by tuning the number of training epochs to stop the training before reaching 100% training accuracy (Carlini et al., 2022). The best hyperparameters are reported at Table˜8.
Table 8: Final hyperparameters used for training the various architecture in Section˜5.
@@ Appendix F Compute resources
All the experiments are run on our internal cluster with the GPU Tesla V100 16GB/32GB of memory. We give an estimate of the amount of compute required for each experiment. For Section˜3, we trained 96 shadow models for Waterbirds and CelebA, and 48 for FMoW, MultiNLI, and CivilComments which took ~600 hours of computing. We trained 16 shadow models for FMoW4 and FMoW16 which took another ~100 hours. For Section˜5, we trained in total 128 shadow models on Waterbirds and CelebA averaging around ~100 hours. Lastly for Section˜4.2, we trained 32 ConvNext-t shadow models and 5 target models for about ~100 hours. The full research required additional computing for hyperparameter grid searches, in particular for differential privacy training which is known to be difficult.